# Milan, Italy

### v3

@@ 5.3 Distinctive Advancements and Limitations
Our research innovates in process simulation by moving beyond the conventional stochastic generative models used in deep learning, focusing instead on an architecture that integrates extensive constraints and a-priori knowledge for a nuanced understanding and simulation of data patterns, diverging from mere sequence prediction. Unlike conventional deep learning solutions [9], which are predominantly stochastic generative models (as elaborated in section 3), our approach takes a step towards overcoming this limitation. Existing methods focus primarily on sequence prediction, which is often similar to, but not synonymous with, simulation. In contrast, a key innovation in CoSMo’s inference stage is its ability to generalize and adaptively manipulate process changes through the constraint vector, represented in this study by $\text{LTL}_{f}$ rules. This feature represents a significant step toward addressing the stochastic nature inherent in deep learning models and incorporating critical information into the learned model to allow more sophisticated evaluations (e.g., what-if analysis). Such capabilities are not only essential for PM practitioners and stakeholders but are also ideal for complex situations where a deep understanding and adherence to an extensive set of rules or a-priori knowledge is essential.
While our study has demonstrated promising outcomes, it is important to acknowledge certain limitations to be addressed in future research. First, the application of the DECLARE language modeling in our case study highlighted some bottlenecks, notably the correlation among some rules (e.g., existence and absence rules), which might introduce redundant information into the learning process. To enhance model efficiency and accuracy, future work will explore preprocessing steps to eliminate these correlated features, thereby streamlining the feature vector and potentially improving performance on complex templates and datasets. Secondly, our focus has been predominantly on methodological validation, with less attention to the architecture’s applicability in real-world settings. The process of rules and activity selections (see Table 2), conducted arbitrarily in this study, would benefit from the involvement of domain experts in practical scenarios to ensure the relevance and impact of chosen rules on the simulations are accurately assessed. Lastly, there is a need to broaden the scope of exploration to include diverse types of constraints and a-priori knowledge, moving beyond declarative rules to incorporate user-based constraints and intra-case feature availability, for instance.
@@ 6 Conclusion
In this work, we introduced a novel recurrent architecture tailored for conditioned process simulation, embodied in the CoSMo framework. Through evaluations of real-world event logs and DECLARE templates, CoSMo demonstrated its ability to adapt to various constraints and datasets, offering process simulations that align with specified conditions. Moreover, CoSMo addresses the inherent stochasticity of deep learning models, enhancing its suitability for process simulation. Acknowledging the discussed limitations, our findings pave the way for future advancements that could significantly improve the application of deep learning-based process simulation models.

### v4

@@ 5.3 Distinctive Advancements and Limitations
Our research innovates in process simulation using deep learning by moving beyond the conventional stochastic generative models, focusing instead on an architecture that integrates extensive constraints and a-priori knowledge for a nuanced understanding and simulation of data patterns, diverging from mere sequence prediction. Unlike conventional deep learning solutions [9], which are predominantly stochastic generative models (as elaborated in section 3), our approach takes a step towards overcoming this limitation. Existing methods focus primarily on sequence prediction, which is often similar to, but not synonymous with, simulation. In contrast, a key innovation in CoSMo’s inference stage is its ability to generalize and adaptively manipulate process changes through the constraint vector, represented in this study by $\text{LTL}_{f}$ rules. This feature represents a significant step toward addressing the stochastic nature inherent in deep learning models and incorporating critical information into the learned model to allow more sophisticated evaluations (e.g., what-if analysis). Such capabilities are not only essential for PM practitioners and stakeholders but are also ideal for complex situations where a deep understanding and adherence to an extensive set of rules or a-priori knowledge is essential.
While our study has demonstrated promising outcomes, it is important to acknowledge certain limitations to be addressed in future research. First, the application of the DECLARE language modeling in our case study highlighted some bottlenecks, notably the correlation among some rules (e.g., existence and absence rules), which might introduce redundant information into the learning process. To enhance model efficiency and accuracy, future work will explore preprocessing steps to eliminate these correlated features, thereby streamlining the feature vector and potentially improving performance on complex templates and datasets. Secondly, our focus has been predominantly on methodological validation, with less attention to the architecture’s applicability in real-world settings. The process of rules and activity selections (see Table 2), conducted arbitrarily in this study, would benefit from the involvement of domain experts for each event log in practical scenarios to ensure the relevance and impact of chosen rules on the simulations are accurately assessed. Lastly, there is a need to broaden the scope of exploration to include diverse types of constraints and a-priori knowledge, moving beyond declarative rules to incorporate, for instance, user-based constraints (such as time-related constraints to change, for instance, the distribution of waiting time) and intra-case feature availability (e.g., resources).
@@ 6 Conclusion
In this work, we introduced a novel recurrent architecture tailored for conditioned process simulation, embodied in the CoSMo framework. Through evaluations of real-world event logs and DECLARE templates, CoSMo demonstrated its ability to adapt to various constraints and datasets, offering process simulations that align with specified conditions. Moreover, CoSMo mitigates the inherent stochasticity of deep learning models, enhancing its suitability for process simulation. Acknowledging the discussed limitations, our findings pave the way for future advancements that could significantly improve the application of deep learning-based process simulation models. Future works include the exploration of additional constraints and a-priori knowledge, practical applications via what-if analysis, the involvement of domain experts for rule selection, and the inclusion of other event attributes to be simulated.